# See or Say Graphs: Agent-Driven Scalable Graph Understanding with Vision-Language Models

### v1

@@ 3.4.1 Visual Graph Thoughts
Existing text-based CoT methods often suffer from pseudo-visual reasoning, in which visual information gradually fades as the reasoning chain lengthens Chen et al. (2024b); Zhang et al. (2024b); Zou et al. (2024); Cheng et al. (2025). To overcome this limitation in multimodal reasoning for complex local tasks, we introduce Visual Graph Thoughts.
@@
$(o_{t},a_{t})=\mathcal{M}_{\text{VRA}}(S_{t})=\mathcal{M}_{\text{VRA}}(G_{\text{image}}^{(t-1)},H_{t-1},{\pi}_{t})$ (2)
If $a_{t}$ invokes $f_{\mathrm{viz}}$ (e.g., node highlighting), the visual representation is updated as $G_{\text{image}}^{(t)}=f_{\mathrm{viz}}(G_{\text{image}}^{(t-1)},a_{t})$; otherwise, $G_{\text{image}}^{(t)}=G_{\text{image}}^{(t-1)}$. This closed-loop process continues until all plan steps are executed.
Finally, $\mathcal{M}_{\text{VRA}}$ aggregates the reasoning trace $H_{n}$ to produce the final answer $A_{\text{final}}=f_{\text{summarize}}(H_{n})$, where $f_{\text{summarize}}$ aggregates intermediate reasoning outputs to generate the final prediction.
@@ 3.4.2 Aligning Visual Graph Thoughts with Process-level DPO
To further improve the multimodal reasoning capability of $\mathcal{M}_{\text{VRA}}$ by promoting the generation of reliable visual graph thoughts, we adopt a process-level Direct Preference Optimization (DPO) training strategy Rafailov et al. (2023). Compared to outcome-based methods Kong et al. (2025a); Liu et al. (2025), we model the entire multimodal reasoning trajectory, which provides stable, step-level reinforcement for graph inference tasks.
@@
We construct a process preference dataset $\mathcal{D}$, where each sample is a triplet $(\mathbf{x},y_{w},y_{l})$. Here, $y_{w}$ denotes the “Chosen” path representing the preferred reasoning trajectory, while $y_{l}$ denotes the “Rejected” path representing a non-preferred trajectory. The construction details are in Appendix A.
@@
, where the policy ratio is defined as $r_{\theta}(y|\mathbf{x})=\pi_{\theta}(y|\mathbf{x})/\pi_{\mathrm{ref}}(y|\mathbf{x})$, and $\beta$ controls the regularization strength. By maximizing the log-probability margin between the preferred ($y_{w}$) and rejected ($y_{l}$) trajectories, the model is guided away from predefined reasoning errors. The resulting policy generates visual graph thoughts that are both logically coherent and tightly grounded in visual graph evidence throughout the inference process.
3.5 Text Modality-based Graph Understanding
For simple property tasks, we adopt GraphRAG-based retrieval for targeted information extraction. Note that existing RAG variants and optimizations Deng et al. (2023); Lee et al. (2025); Gao et al. (2025); Chang et al. (2025) are orthogonal to GraphVista and can be easily integrated to improve retrieval quality. To maintain generality, we employ the standard RAG formulation in this work.
For a given task request $(Q,\mathcal{E})$, we first employs a retrieval function $f_{\text{retrieve}}$ to question $\mathcal{K}$ and extract relevant context, formalized as $C_{\text{retrieved}}=f_{\text{retrieve}}(\mathcal{K},(Q,\mathcal{E}))$. Subsequently, the downstream VLM is prompted with both the original question $Q$ and the retrieved context $C_{\text{retrieved}}$ to generate the final answer.
Table 1: Performance comparison on Grena and GraphSQA Benchmark. Red and Blue highlight the best and second-best results respectively.
@@ 3.5 Text Modality-based Graph Understanding
Simple Complex Overall Simple Complex Overall
Text-only InternVL3-9B 0.1211 0.0287 0.0829 0.4788 0.2498 0.3169
GLM-4.1V-9B 0.0157 0.0071 0.0121 0.2532 0.1670 0.1929
Qwen2.5-VL-7B 0.0096 0.0095 0.0096 0.2340 0.1122 0.1488
Qwen3-8B 0.0065 0.0056 0.0061 0.1351 0.1352 0.1352
Gemma-3-12B 0.0369 0.0358 0.0364 0.5917 0.3245 0.4033
GraphPRM (DPO) Qwen2.5-VL-7B (DPO) 0.0332 0.0087 0.0231 0.7198 0.3455 0.4550
Qwen3-8B (DPO) 0.0060 0.0049 0.0055 0.3629 0.2357 0.2728
Gemma-3-12B (DPO) 0.0417 0.0335 0.0383 0.6235 0.3455 0.4195
GraphInsight InternVL3-9B 0.1788 0.0056 0.1073 0.7564 0.2315 0.3860
GLM-4.1V-9B 0.1849 0.0016 0.1092 0.4313 0.3273 0.3583
Qwen2.5-VL-7B 0.1925 0.0085 0.1165 0.5769 0.2293 0.3348
Gemma-3-12B 0.2919 0.0236 0.1811 0.7372 0.2478 0.3779
GraphToken Qwen2.5-VL-7B 0.0981 0.0782 0.0899 0.0388 0.1770 0.1376
Qwen3-8B 0.1589 0.0757 0.1245 0.1742 0.1508 0.1579
Gemma-3-12B 0.0158 0.0576 0.0331 0.0566 0.1091 0.0927
GITA InternVL3-9B 0.1751 0.1098 0.1481 0.3910 0.1755 0.2402
GLM-4.1V-9B 0.1651 0.1175 0.1454 0.3397 0.1562 0.2112
Qwen2.5-VL-7B 0.1351 0.0632 0.1054 0.3205 0.1596 0.2079
Gemma-3-12B 0.1732 0.1435 0.1609 0.2692 0.1739 0.2026
GraphVista (Ours) InternVL3-9B 0.9357 0.3578 0.6970 0.7970 0.6886 0.7211
GLM-4.1V-9B 0.9387 0.2727 0.6637 0.7954 0.6657 0.7046
Gemma-3-12B 0.9362 0.3477 0.6932 0.7991 0.6815 0.7168
Gemma-3-4B (DPO) 0.9362 0.3397 0.6899 0.7943 0.6780 0.7129
Qwen2.5-VL-7B (DPO) 0.9367 0.3912 0.7114 0.7973 0.6806 0.7156
4 Evaluation
@@ 4.1 Experimental Setup
4.1.1 Baselines
We compare GraphVista with representative state-of-the-art baselines across three categories: (a) Text-based methods, which reason solely over textual graph descriptions, including the standard Text-only approach, GraphPRM Peng et al. (2025) trained via DPO, and GraphInsight Cao et al. (2025). (b) GNN-based methods, which encode structural information into latent representations using graph neural networks. A representative example is GraphToken [^5] These methods are orthogonal to GraphVista, focusing on graph encoding rather than improving the intrinsic graph understanding of VLMs. We include them for completeness. Perozzi et al. (2024). (c) Hybrid methods, exemplified by GITA Wei et al. (2024), which integrate both visual and textual graph modalities into VLMs.
4.1.2 Benchmarks and Evaluation Tasks
Existing graph understanding benchmarks suffer from limited coverage of graph scales and topologies, a lack of task categorization grounded in fundamental graph properties, and insufficient support for evaluating multi-step visual reasoning. To address these gaps, we introduce Grena, a large-scale, multi-task benchmark we designed to evaluate graph understanding of LLMs and VLMs, which encompasses a wide range of task types found in existing benchmarks Wang et al. (2023); Yuan et al. (2025a); Zhu et al. (2025).
Grena spans a wide diversity of graph scales, with node counts ranging up to 2050, and encompasses 18 distinct task types, which are organized into two categories: (1) 7 Simple Property Tasks (Simple) focus on querying basic properties and local structures, and (2) 11 Complex Local Tasks (Complex) require fine-grained topological information and performing fine-grained inference. Further details are provided in Appendix A.
In our experiments, we use Grena to evaluate graph understanding performance on large-scale graphs ($|V|\in[50,2050]$), covering a total of 13,860 tasks, which are split into 8,136 simple and 5,724 complex instances. Complementarily, we use the GraphSQA benchmark Cao et al. (2025) for small-graph understanding, with $|V|$ in the range of $[15,50]$. Across all tasks and benchmarks, we use Accuracy (ACC) as the evaluation metric. Methods that require DPO are trained on $\mathcal{D}$.
(a) Complex (Gemma-3)
(b) Simple (Gemma-3)
(c) Complex (Qwen2.5-VL)
(d) Simple (Qwen2.5-VL)
@@ 4.1.2 Benchmarks and Evaluation Tasks
Table 2: Ablation study on reasoning strategies of GraphVista. The best results within each VLM group are bolded, and the second-best results are underlined.
Strategy Simple Complex Overall
InternVL3-9B
CoT (Baseline) 0.9368 0.1422 0.6086
Visual Graph Thoughts (Text) 0.2220 0.6416
Visual Graph Thoughts (Visual) 0.3578 0.6977
Gemma-3-12B
CoT (Baseline) 0.9362 0.1406 0.6076
Visual Graph Thoughts (Text) 0.2447 0.6506
Visual Graph Thoughts (Visual) 0.3477 0.6932
GLM-4.1V-9B
CoT (Baseline) 0.9387 0.1406 0.6091
Visual Graph Thoughts (Text) 0.2300 0.6460
Visual Graph Thoughts (Visual) 0.2727 0.6637
Qwen2.5-VL-7B (DPO)
CoT (Baseline) 0.9367 0.1414 0.6083
Visual Graph Thoughts (Text) 0.3429 0.6915
Visual Graph Thoughts (Visual) 0.3912 0.7114
4.1.3 Models
We instantiate GraphVista and all baselines using recent open-source, competitive VLMs, including Gemma-3-12B, Gemma-3-4B Kamath et al. (2025), Qwen2.5-VL-7B Bai et al. (2025), InternVL3-9B Zhu and et al (2025), and GLM-4.1V-9B-Thinking Hong et al. (2025), a reasoning-oriented VLM. For text-based methods, we include more capable LLMs such as Qwen3-8BYang et al. (2025) to ensure a fairer comparison.
@@ 4.2 Performance
As shown in Table 1, GraphVista significantly outperforms all baseline methods on both Grena and GraphSQA benchmarks.
4.2.1 Simple Property Tasks
@@
4.2.2 Complex Local Tasks
@@
Table 3: Ablation results comparing frozen base models (“Frozen”) with models further optimized via DPO (“DPO”).Red and Blue highlight the best and second-best results respectively.
Model Strategy Simple Complex Overall
Gemma-3-4B Frozen 0.9362 0.2933 0.6707
DPO 0.3397 0.6899
Qwen2.5-VL-7B Frozen 0.9367 0.3240 0.6837
DPO 0.3912 0.7114
@@ 4.3 Ablation Study
4.3.1 Effectiveness of Reasoning Strategies
We evaluate the effect of visual grounding by comparing Visual Graph Thoughts (Visual) with its text-only variant and the standard textual CoT, where visual subgraphs are replaced by textual descriptions. In Table 2, the visual variant consistently outperforms the baselines, especially on complex tasks, achieving up to $1.1\times$ and $2.8\times$ higher accuracy than the text-only and CoT settings, respectively. These highlight the importance of explicit structural grounding, as purely textual reasoning often yields ambiguous interpretations.
4.3.2 Impact of DPO
We evaluate the impact of process-level DPO by comparing fine-tuned models with their frozen-base counterparts. As reported in Table 3, DPO enhances performance across all settings, with the most significant gains on complex tasks. For instance, DPO boosts the accuracy of Qwen2.5-VL-7B on complex tasks from 0.3240 to 0.3912. These results validate that aligning model generation with expert visual reasoning trajectories is highly effective for complex local tasks.
(a) Gemma-3 (b) Qwen2.5-VL
Figure 4: Hyperparameter analysis on $k$ and $N_{max}$
Figure 5: Hyperparameter analysis on storage strategy of $\mathcal{K}$
@@ 4.4 Hyperparameter Analysis
4.4.1 Subgraph Extraction Parameters $k$ and $N_{max}$
We analyze the impact of $k$ and $N_{max}$ on performance for complex local tasks. As shown in Figure 4, Gemma-3 reaches its highest accuracy with a 2-hop neighborhood, indicating that it can leverage a broader structural context. In contrast, Qwen2.5-VL performs best with a 1-hop neighborhood, suggesting that larger subgraphs may introduce extraneous information. These observations point to the need for adapting subgraph extraction strategies to the characteristics of different VLM architectures.
4.4.2 Storage Strategy of $\mathcal{K}$
We analyze the effect of pruning node tiers in $\mathcal{K}$, which are defined by the parameters $K_{1}$% and $K_{2}$%. As demonstrated in Figure 5, performance on simple property tasks improves as more tiers of the graph are stored. Retaining just the top $K_{1}$% of nodes (Tier 1) along with their 2-hop neighborhoods already provides a strong baseline for localized retrieval. Adding the next $K_{2}$% of nodes (Tier 2) enhances performance further, although the gains diminish when including peripheral nodes (Tier 3). While full graph storage is essential for precise macroscopic tasks where any pruning leads to information loss, these findings reveal a clear trade-off between performance and storage efficiency for general reasoning tasks.

### v2

@@ 3.4.1 Visual Graph Thoughts
Existing text-based CoT methods often suffer from pseudo-visual reasoning, in which visual information gradually fades as the reasoning chain lengthens Chen et al. (2024b); Zhang et al. (2024b); Zou et al. (2024); Cheng et al. (2025). To overcome this limitation in multimodal reasoning, we introduce Visual Graph Thoughts.
@@
$(o_{t},a_{t})=\mathcal{M}_{\text{VRA}}(S_{t})=\mathcal{M}_{\text{VRA}}(G_{\text{image}}^{(t-1)},H_{t-1},{\pi}_{t}).$ (2)
If $a_{t}$ invokes $f_{\mathrm{viz}}$ (e.g., node highlighting), the visual representation is updated as $G_{\text{image}}^{(t)}=f_{\mathrm{viz}}(G_{\text{image}}^{(t-1)},a_{t})$; otherwise, $G_{\text{image}}^{(t)}=G_{\text{image}}^{(t-1)}$. This closed-loop process continues until all plan steps are executed. Finally, $\mathcal{M}_{\text{VRA}}$ aggregates the reasoning trace $H_{n}$ to produce the final answer.
@@ 3.4.2 Aligning Visual Graph Thoughts with Process-level DPO
To further improve the multimodal reasoning capability of $\mathcal{M}_{\text{VRA}}$ by promoting the generation of reliable visual graph thoughts, we adopt the process-level Direct Preference Optimization (DPO) training strategy Rafailov et al. (2023), which models the full multimodal reasoning trajectory and provides stable, step-level supervision for graph inference. Advanced reinforcement learning methods Hu et al. (2025); Dinucu-Jianu et al. (2025); Yu et al. (2025) are orthogonal to our framework and can be naturally integrated. We adopt the standard process-level DPO formulation to ensure generality and training stability.
@@
We construct a process preference dataset $\mathcal{D}$, where each sample is a triplet $(\mathbf{x},y_{w},y_{l})$. Here, $\mathbf{x}$ denotes the multimodal input, $y_{w}$ represents the “Chosen” path (preferred reasoning trajectory), and $y_{l}$ represents the “Rejected” path (non-preferred trajectory). The construction details are in Appendix A.
@@
, where the policy ratio is defined as $r_{\theta}(y|\mathbf{x})=\pi_{\theta}(y|\mathbf{x})/\pi_{\mathrm{ref}}(y|\mathbf{x})$, and $\beta$ controls the regularization strength. Maximizing the log-probability margin between preferred ($y_{w}$) and rejected ($y_{l}$) trajectories effectively suppresses predefined reasoning errors. The resulting policy generates visual graph thoughts that are logically coherent and faithfully grounded in visual evidence. Training details are available in Appendices E and F.
3.5 Text Modality-based Graph Structure Understanding
For text-modality tasks, we adopt GraphRAG-based retrieval for targeted information extraction. Note that existing RAG variants and optimizations Deng et al. (2023); Lee et al. (2025); Gao et al. (2025); Chang et al. (2025) are orthogonal to GraphVista and can be easily integrated to improve retrieval quality. To maintain generality, we employ the standard RAG formulation in this work.
Given the parsed tuple $(Q,\mathcal{E})$, the relevant context $\mathcal{C}$ is extracted from $\mathcal{K}$ by aggregating the structural descriptions of entities in $\mathcal{E}$. The VLM then generates the answer conditioned on both $Q$ and $\mathcal{C}$.
4 Evaluation
Table 1: Performance comparison on Grena and GraphSQA Benchmark. Red and Blue highlight the best and second-best results, respectively.
@@ 4 Evaluation
Text Visual Collab. Overall Text Visual Overall
Text-only GLM-4.1V-9B 0.0352 0.0055 0.0110 0.0217 0.2532 0.1670 0.1929
InternVL3-9B 0.0452 0.0180 0.0303 0.0345 0.4788 0.2498 0.3185
Qwen2.5-VL-7B 0.0558 0.0345 0.0337 0.0450 0.2340 0.1122 0.1487
Qwen3-8B 0.0403 0.0367 0.0096 0.0321 0.1351 0.1352 0.1352
Gemma-3-12B 0.1246 0.0591 0.0337 0.0858 0.5917 0.3245 0.4047
GPT-5-mini 0.1451 0.0933 0.0515 0.1093 0.6464 0.6082 0.6197
GraphPRM (DPO) Qwen2.5-VL-7B (DPO) 0.0589 0.0410 0.0314 0.0477 0.7198 0.3455 0.4578
Qwen3-8B (DPO) 0.0446 0.0379 0.0068 0.0339 0.3629 0.2357 0.2739
Gemma-3-12B (DPO) 0.1281 0.0590 0.0434 0.0899 0.6235 0.3455 0.4289
GraphInsight GLM-4.1V-9B 0.1951 0.2655 0.2172 0.2189 0.4313 0.3273 0.3585
InternVL3-9B 0.2507 0.2915 0.2088 0.2515 0.7564 0.2315 0.3890
Qwen2.5-VL-7B 0.2011 0.1078 0.0976 0.1520 0.5769 0.2293 0.3336
Gemma-3-12B 0.2469 0.2357 0.1131 0.2123 0.7372 0.2478 0.3946
GPT-5-mini 0.3031 0.3429 0.2843 0.3091 0.7211 0.6446 0.6676
GraphToken Qwen2.5-VL-7B 0.0141 0.0982 0.0396 0.0423 0.0388 0.1770 0.1355
Qwen3-8B 0.0052 0.0665 0.0089 0.0222 0.1742 0.1508 0.1578
Gemma-3-12B 0.0191 0.1064 0.0475 0.0488 0.0566 0.1091 0.0933
GITA GLM-4.1V-9B 0.0665 0.2152 0.1835 0.1333 0.3397 0.1562 0.2112
InternVL3-9B 0.0714 0.1045 0.1546 0.0998 0.3910 0.1755 0.2401
Qwen2.5-VL-7B 0.1099 0.1799 0.2271 0.1561 0.3205 0.1596 0.2079
Gemma-3-12B 0.1218 0.1591 0.2344 0.1583 0.2692 0.1739 0.2025
GraphVista (Ours) GLM-4.1V-9B 0.9352 0.3540 0.2559 0.6214 0.7954 0.6657 0.7046
InternVL3-9B 0.9357 0.3701 0.3449 0.6469 0.7970 0.6886 0.7211
Qwen2.5-VL-7B (DPO) 0.9363 0.4309 0.4478 0.6876 0.7973 0.6806 0.7156
Gemma-3-12B (DPO) 0.9361 0.4431 0.6018 0.7272 0.7991 0.6815 0.7168
GPT-5-mini 0.9406 0.4526 0.6241 0.7372 0.8497 0.7967 0.8126
@@ 4.1 Experimental Setup
Baselines.
We compare GraphVista with state-of-the-art baselines across three categories: (a) Text-based methods, which reason exclusively over textual descriptions, including the standard Text-only baseline, GraphPRM Peng et al. (2025) (trained via DPO), and GraphInsight Cao et al. (2025). (b) GNN-based methods, such as GraphToken Perozzi et al. (2024), which encode structural information into latent representations [^9] These methods focus on graph encoding rather than VLM intrinsic understanding. We include them for completeness.. (c) Hybrid methods, represented by GITA Wei et al. (2024), which integrate both visual and textual modalities.
Benchmarks and Evaluation Metrics.
Current graph understanding benchmarks suffer from restricted scale and topological diversity, with insufficient support for multi-step visual reasoning. We introduce Grena, a large-scale benchmark designed to evaluate VLMs across 20 distinct task types, covering the scope of existing baselines Wang et al. (2023); Yuan et al. (2025a); Zhu et al. (2025b). Grena spans a wide range of graph scales (up to 2,050 nodes) and topologies. Further details are provided in Appendix A.
We use Grena to evaluate graph structure understanding performance on large-scale graphs ($|V|\in[50,2050]$), covering a total of 22,800 tasks. Complementarily, we use the GraphSQA benchmark Cao et al. (2025) for small-graph structure understanding, with $|V|$ in the range of [15,50][15,50] [^10] As this small-graph benchmark lacks modality-collaborative tasks, we do not separate local and global tasks. Instead, we directly adopt its original classification, referring to the two categories as “Text” and “Visual”.. Across all tasks and benchmarks, we use Accuracy (ACC) as the evaluation metric. Methods requiring DPO are trained on $\mathcal{D}$, as detailed in Appendix E.
Table 2: Ablation study of reasoning strategies in GraphVista. Best results within each VLM group are bolded, and second-best results are underlined.
Strategy Text Visual Collab. Overall
GLM-4.1V-9B
CoT (Baseline) 0.9352 0.1376 0.1667 0.5433
Visual Graph Thoughts (Text) 0.2734 0.1902 0.5846
Visual Graph Thoughts (Visual) 0.3540 0.2559 0.6214
InternVL3-9B
CoT (Baseline) 0.9357 0.1422 0.1332 0.5368
Visual Graph Thoughts (Text) 0.2879 0.2511 0.6031
Visual Graph Thoughts (Visual) 0.3701 0.3449 0.6469
Qwen2.5-VL-7B (DPO)
CoT (Baseline) 0.9363 0.1691 0.1684 0.5525
Visual Graph Thoughts (Text) 0.3629 0.3355 0.6431
Visual Graph Thoughts (Visual) 0.4309 0.4478 0.6876
Gemma-3-12B (DPO)
CoT (Baseline) 0.9361 0.1676 0.1703 0.5525
Visual Graph Thoughts (Text) 0.3551 0.4200 0.6610
Visual Graph Thoughts (Visual) 0.4431 0.6018 0.7272
Models.
We instantiate all methods using recent open-source, competitive VLMs, including Gemma-3-12B Kamath et al. (2025), Qwen2.5-VL-7B Bai et al. (2025), InternVL3-9B Zhu et al. (2025a), and GLM-4.1V-9B-Thinking Hong et al. (2025), a reasoning-oriented VLM. For text-based methods, we include more capable LLMs such as Qwen3-8B Yang et al. (2025) to ensure a fairer comparison. For training-free methods, we include GPT-5-mini to assess generality.
Table 3: Ablation comparing frozen base models (“Frozen”) and DPO-fine-tuned models (“DPO”). Red and Blue denote the best and second-best results.
Model Strategy Text Visual Collab. Overall
Qwen2.5-VL-7B Frozen 0.9363 0.3951 0.3589 0.6571
DPO 0.4309 0.4478 0.6876
Gemma-3-12B Frozen 0.9361 0.4005 0.5387 0.7010
DPO 0.4431 0.6018 0.7272
(a) Text (Gemma-3)
(b) Text (Qwen2.5-VL)
(c) Visual (Gemma-3)
(d) Visual (Qwen2.5-VL)
(e) Collab. (Gemma-3)
(f) Collab. (Qwen2.5-VL)
@@ 4.2 Performance
As shown in Table 1, GraphVista consistently outperforms all baselines on both benchmarks.
Text-modality Tasks.
@@
Visual and Modality-Collaborative Tasks.
@@ 4.3 Ablation Study
Reasoning Strategies.
[…]
Impact of DPO.
Comparing fine-tuned models with frozen baselines (Table 3) demonstrates that process-level DPO enhances performance across all settings, particularly for complex tasks. Notably, DPO improves Qwen2.5-VL-7B’s accuracy on collaborative tasks from 0.3589 to 0.4478. This verifies the effectiveness of aligning model outputs with expert visual reasoning trajectories.
(a) Gemma-3 (Visual)
(b) Gemma-3 (Collab.)
(c) Qwen2.5-VL (Visual)
(d) Qwen2.5-VL (Collab.)
Figure 4: Hyperparameter analysis on $k$ and $N_{\max}$
Figure 5: Tier number analysis in $\mathcal{K}$
@@ 4.4 Hyperparameter Analysis
Subgraph Extraction.
We analyze the impact of $k$ and $N_{max}$ on performance for visual-modality tasks. As shown in Figure 4, Gemma-3 reaches its highest accuracy with a 2-hop neighborhood, indicating that it can leverage a broader structural context. In contrast, Qwen2.5-VL performs best with a 1-hop neighborhood, suggesting that larger subgraphs may introduce extraneous information. These observations point to the need for adapting subgraph extraction strategies to the characteristics of different VLM architectures.
Number of Tiers in $\mathcal{K}$.
We examine the impact of storage granularity in $\mathcal{K}$ across three settings: Tier 1 (Core), Tier 1+2 (Backbone), and Full Graph. As shown in Figure 5, performance positively scales with structural coverage. Tiers 1 and 2 provide efficient baselines, but adding Tier 3 significantly improves performance (e.g., 0.936 on InternVL3-9B) at the cost of higher storage. This underscores a clear trade-off: prioritizing central nodes ensures efficiency, whereas full graph coverage is essential for maximizing accuracy.